The data for a certain test is provided below





X 100 200 300 400 500 600 700




Y 30 50 60 80 100 110 130





Compute the value of the correlation.




Find the equation of the regression lime.


1
Expert's answer
2021-12-06T17:38:32-0500

 correlation:

"r=\frac{\sum (x_i-\overline{x})(y_i-\overline{y})}{\sqrt{\sum (x_i-\overline{x})^2\sum(y_i-\overline{y})^2}}=0.9972"


regression line:

"y=ax+b"


"a=\frac{\sum (x_i-\overline{x})(y_i-\overline{y})}{\sum (x_i-\overline{x})^2}=0.1643"


"b=\overline{y}-a\overline{x}=14.2857"


"y=0.1643x+14.2857"